Twenty students ploay basketball and 14 students play football. If there are 30 students all together, how many both basket ball and football?
Answer by venugopalramana(3286)

Twenty students ploay basketball and 14 students play football. If there are 30 students all together, how many both basket ball and football?
You may edit the question. Maybe convert formulae to the same formula notation
total playing basket ball & foot ball=20+14=34
but there are only 30 students.hence assuming that there are none who do not play either game ,there must be 34-30=4 playing both foot ball and basket ball